OpenStudy (anonymous):

Solve the following system of equations for y: 5x + 4y = 61 2x + 4y = 70 A. y = 17 B. None of these. C. y = -3 D. y = 15 E. y = 19

OpenStudy (anonymous):

Subtract the second equation from the first one, this will eliminate y, then find x. Put that x back in any of the equations, this gives you y.

OpenStudy (anonymous):

5x + 4y = 61 --(1) 2x + 4y = 70 --(2) (1)-(2): 3x=-9 x=-3 sub this into (2) 2(-3)+4y=70 -6+4y=70 y=19

OpenStudy (anonymous):

E